### v2

# From shape to fate: making bacterial swarming expansion predictable

Abstract
Microbial swarming on mucosal surfaces reshapes microbial communities and influences mucosal healing and antibiotic tolerance. Yet even with time-lapse microscopy and deep learning, analyses of swarming colonies remain largely descriptive and lack accurate forecasting of front reorganization. This limitation is significant because the advancing edge determines access to nutrients, host tissues, and competing microbes. The expansion of Enterobacter sp. SM3 swarms is recast as a problem of morphological forecasting, and SwarmEvo is assembled as a time-lapse dataset of boundary-resolved segmentations. TexPol–Net, a texture- and geometry-aware segmentation model, sharpens diffuse edges and preserves finger-like fronts, creating a stable substrate for dynamics. On this representation, Morpher is developed as an autoregressive forecasting network with a “Morphon” memory that links local curvature to long-range temporal dependencies. Morpher outperforms leading video-prediction models in maintaining front localization and anisotropic branching, and modest segmentation improvements yield noticeably more stable forecasts. Ablations across sequence models, inference strategies, and observation ratios show that attention-based architectures with structural memory best preserve branching propagation dynamics. By uniting geometry-aware segmentation with morphology-level forecasting, this framework turns swarming expansion into a predictive dynamical system, enabling quantitative interrogation and potential control of microbial collectives during mucosal repair and gut ecosystem engineering.
Keywords: bacterial swarming; morphological forecasting; collective behavior; spatiotemporal dynamics; deep learning; predictive modeling
Significance Statement Swarming bacteria reshape their environment through a moving front that governs access to oxygen, host tissue, and competing microbes. Yet this front has largely been described rather than predicted. We show that swarming trajectories are not organized by experimental conditions but by colony-specific dynamics, shifting prediction away from condition labels toward direct measurement of the front itself. By recasting the front as a geometric state whose evolution can be forecast, our framework links measurement fidelity to predictive stability. This establishes colony morphology as a quantitative, forward-looking variable and opens a path toward anticipating—and ultimately controlling—collective microbial behavior.
1 Introduction
Microbial communities form dense and continually reorganizing ecological networks within and around animal hosts [40, 10, 4, 69, 55, 53]. Their spatial organization and collective movement shape population expansion, interspecies interactions, and tissue homeostasis in both health and disease [23, 39, 33, 20]. Among these behaviors, collective surface motility on semi-solid substrates—classically termed swarming [1, 8, 66, 29, 3, 6]—produces continually reconfiguring fronts and colony architectures that both reflect coordinated cellular behavior and actively remodel the local host environment [14, 68, 45, 25]. Although classic studies of swarm pattern formation, cell-state differentiation, and periodic colony expansion documented rich phenomenology [50, 49, 24, 28], they stopped short of treating front evolution as a predictive problem. We still lack a framework that can predict how a swarming colony will change shape over time, a capability that would turn morphology from a visual endpoint into a quantitative variable for probing how microbial collectives respond to environmental and physiological cues.
This gap persists despite substantial progress in colony analysis. Early studies were limited by sparse imaging and qualitative interpretation [35, 41]. Classical computer-vision pipelines improved colony detection and counting [7, 12, 54, 71, 2, 70], but they deteriorate when boundaries blur, textures reorganize, or colonies overlap. Physical analyses of active suspensions and deformable colony interfaces have revealed turbulence-like flows and curvature-dependent edge dynamics [26, 65], yet they often assume approximate symmetry or compress growth into low-dimensional summaries that cannot capture the irregular, anisotropic, and burst-like propagation of swarming fronts. Deep learning has broadened biological image analysis [72, 37, 19, 48] through CNN-based colony detectors [17, 64], temporal classifiers of motility states [44], and hybrid detection-growth pipelines [42]. Advances in imaging, including coherent time-lapse microscopy for early species identification [60] and engineered swarming biosensors [15], further emphasize how much biological information is encoded in colony morphology. But these approaches remain fundamentally descriptive: they detect, classify, or summarize; they do not forecast contour evolution. Even single-frame predictors of motility type [34] collapse a dynamic process into a static label.
A central reason is that swarming violates the appearance continuity assumed by most natural-scene video models. Fronts advance through intermittent bursts, transient asymmetries, and rapid multiscale reorganization. Curvature modulates local speed; protrusions emerge and retract discontinuously; texture shifts without preserving pixel-level coherence. Under these conditions, extrapolating appearance is neither stable nor biologically persuasive. What carries the future is the front itself: its geometry, branching structure, and temporal continuity. In spatially expanding microbial populations, the advancing frontier constitutes a thin, dynamically active region that governs large-scale structure while the interior remains effectively frozen [21]. A useful predictive model must therefore resolve colony-specific morphology rather than rely on coarse labels, bulk summaries, or framewise appearance alone. Forecasting future shape in this setting is not an image-generation problem but a problem of front dynamics.
This challenge is especially consequential in Enterobacter sp. SM3, a representative swarming commensal from the murine gut [47, 11, 27]. SM3 reshapes intestinal microbial organization and promotes mucosal healing, whereas swarming-deficient mutants lose these beneficial effects [14]. More broadly, swarming has been linked to antibiotic tolerance, virulence regulation, and robust colonization of host-associated surfaces [32, 43, 46], and its dynamics are strongly modulated by surface biochemical cues such as mucin [45]. In inflamed intestinal environments, the advancing swarming front is the actionable interface: its future position determines where oxygen is depleted and anaerobic niches emerge, thereby shaping downstream community reorganization and recovery (Figure 1). Anticipating that front in advance would provide a principled basis for spatially and temporally targeted intervention. Yet no framework currently forecasts swarming morphology at the resolution of individual fronts, leaving a gap between microscopic motility programs and macroscopic pattern evolution.
Here we address that gap by recasting swarming colony expansion as a problem of morphological forecasting in a geometric state space. We assemble the Swarming Morphogenesis Evolution (SwarmEvo) dataset, a time-lapse resource of Enterobacter sp. SM3 expanding across systematically varied semi-solid environments. We then recover boundary-resolved colony states with TexPol–Net, designed to preserve diffuse but biologically important front structures, and forecast their evolution with Morpher, a spatiotemporal model that treats prediction as contour dynamics rather than future-image synthesis. By linking boundary measurement to long-horizon front forecasting, this framework reframes swarming as a measurable and predictable dynamical process, opening a route toward quantitative interrogation—and ultimately control—of microbial collective behavior in living environments.
2 Results
2.1 A state-based formulation of swarming morphology
Swarming expansion in Enterobacter sp. SM3 spans a reproducible morphological spectrum. Even under the same experimental conditions, colonies form two distinct morphogenetic regimes: an anisotropic branching regime (finger-like fronts) and a near-concentric regime (approximately isotropic expansion) (Figure 1). These regimes reflect structured morphogenetic states rather than frame-level visual fluctuations.
We therefore formulate swarming as a two-stage problem: first, measuring the advancing front as a geometric state; second, forecasting the evolution of that state. Time-lapse images are converted into boundary-resolved masks that preserve protrusions while suppressing appearance variability, yielding a stable representation across time. Forecasting is then posed as front evolution rather than future-image synthesis. This shift moves the problem from appearance prediction to state evolution, with morphology serving as a compact state descriptor of colony dynamics.
Figure 1: From swarming dynamics to predictive guidance. In colitis, Enterobacter sp. SM3 swarms along the inflamed mucosal surface, where the advancing front governs access to oxygen and microbial competition. Anticipating its future position enables spatially and temporally targeted intervention. Time-lapse assays are converted by TexPol–Net into boundary-resolved colony states, and Morpher forecasts their evolution. This framework links the biological problem to a two-stage formulation of boundary measurement and morphology-level prediction.
2.2 Individual variability in swarming dynamics
Before evaluating forecasting models, we asked whether swarming trajectories are cleanly organized by assay conditions. They are not. In PCA of trajectory-level perimeter and area features, colonies measured under different temperatures, humidities, and agar concentrations remain extensively overlapped rather than forming distinct clusters (Figure 2a,b). The same pattern persists at the distribution level: for both perimeter and area trajectories, within- and across-condition pairwise distances nearly coincide, with overlap coefficients remaining high (0.92–0.97) and Cliff’s $\delta$ close to zero across all three variable groupings (Figure 2c,d).
Figure 2: Large individual variability and weak separability in swarming morphology. a,b PCA of trajectory-level perimeter and area features from 81 colonies across temperature, humidity, and agar concentration. Broad overlap indicates weak separability by nominal condition. c,d Pairwise-distance distributions for perimeter and area trajectories under the same groupings. Overlap between within- and across-condition distances shows that individual variation is comparable to condition-level differences.
Swarming trajectories are not organized by nominal experimental conditions. Instead, variation is dominated by colony-specific dynamics, with within- and across-condition distances remaining statistically indistinguishable. This eliminates condition labels as a predictive axis and shifts the problem to geometry-resolved state evolution.
2.3 Segmentation defines the predictive state
Because forecasting depends on front geometry, we first evaluated how accurately different segmentation backbones reconstruct colony boundaries across representative growth regimes (Figure 3b). The clearest separation appears in the anisotropic branching regime, where boundary fidelity is most demanding. YOLOv11 captures the global outline but shortens or fragments slender fingers. SAM and SAM2 preserve the colony core while suppressing distal protrusions, driving the front toward a smoother, more circular contour. TexPol–Net remains aligned with both the outer envelope and fine branches, with YOLOv12 as the closest competitor.
Figure 3: TexPol–Net improves colony-front segmentation by coupling texture-sensitive boundary encoding with a geometry-aligned context prior. a TexPol–Net architecture within a prototype-based instance segmentation pipeline. TEA in the backbone preserves boundary texture, and PCA in the bidirectional neck maintains polar consistency during multi-scale fusion. b Qualitative comparison on representative anisotropic branching and near-concentric regimes. TexPol–Net better preserves branch heterogeneity and boundary position than YOLOv11 [30], SAM [31], and SAM2 [51], with YOLOv12 [58] as the closest competitor. c AP as a function of IoU threshold on the SwarmEvo segmentation test set (78 colonies), highlighting stronger performance under boundary-stringent matching. d Image-wise IoU distribution on the same test set, summarizing accuracy across samples. e Image-wise Dice distribution on the same test set, reflecting consistency and variability of segmentation quality.
[…]
2.4 Limits of appearance-based prediction
Morphology-level forecasting is ultimately judged by whether the active front is localized and whether its fine branches are preserved. Region overlap (mIoU) can remain high while the contour drifts; we therefore treat boundary-sensitive distances (HD95 and ASSD) as the primary indicators of predictive fidelity, and use overlap as a complementary check on coarse extent (Figure 4a).
We compared Morpher with representative video prediction models, including MAU [9], MIM [63], PredRNN [61], PredRNNv2 [62], the original SimVP implementation with the TAU temporal unit [18, 57], and the improved SimVPv2 variant with the gSTA module [56], all retrained under identical splits and evaluated on the same 80% observation / 20% prediction protocol (Figure 4a). This ordering holds across all metrics. MIM and SimVP+gSTA preserve coarse regional extent to some degree, reaching mIoU values of 89.32% and 90.52%, but remain substantially worse in boundary accuracy. PredRNN and PredRNNv2 exhibit still larger HD95 and ASSD, indicating stronger drift during extrapolation. SimVP+TAU remains intermediate without resolving the overlap–boundary tradeoff.
Figure 4: Failure modes of generic video prediction in swarming morphology forecasting and the role of morphology-aware representation. a Performance under an 80% observation / 20% prediction protocol. Morpher achieves the best overall accuracy (95.42% mIoU, 10.61 px HD95, 3.93 px ASSD), indicating improved front localization and boundary fidelity. b Long-horizon forecasts on two representative sequences. Generic predictors smooth finger-like branches or accumulate boundary drift, whereas Morpher maintains a coherent front closer to the ground truth across regimes. c Effect of segmentation backbone on forecasting. TexPol–Net masks yield more accurate final-frame predictions than YOLOv12 [58], increasing IoU from 94.21% to 96.63% in the anisotropic branching regime and from 93.30% to 96.38% in the near-concentric regime.
Morpher performs best on all three metrics, reaching 95.42% mIoU, 10.61 px HD95, and 3.93 px ASSD. Relative to the strongest baseline, SimVP+gSTA, this corresponds to a 5.4% gain in overlap and reductions of 42.0% and 55.7% in HD95 and ASSD. These gains are obtained under an 80% observation / 20% prediction protocol, where prediction is restricted to the late expansion stage, when the front is already extended and highly sensitive to small geometric deviations.
The qualitative failures follow the same ordering (Figure 4b). In the anisotropic branching regime, MIM progressively smooths protrusions; SimVP variants retain a coarse outline but truncate lobe tips and blunt high-curvature sectors; MAU, PredRNN, and PredRNNv2 lag in propagation, yielding fronts that are spatially plausible but temporally delayed. In the late near-concentric regime, the same models underestimate radial extent or accumulate boundary drift. Morpher remains closest to the true front across both regimes, preserving local perturbations without losing global scale. The separation reflects a mismatch in what is being modeled: generic video predictors favor appearance continuity, whereas Morpher treats morphology itself as the evolving state.
2.5 Segmentation errors accumulate during forecasting
A causal link between measurement fidelity and forecast stability can be tested by a controlled swap: holding the temporal model fixed while changing only the segmentation backbone that defines the state. We therefore paired the same Morpher architecture with either TexPol–Net or YOLOv12. The static difference is small: TexPol–Net achieves 92.48% mAP50:95, whereas YOLOv12 reaches 91.81%. The temporal consequence is not. Over the final 20% prediction window, YOLOv12 masks incur a 2.4–3.1 IoU-point loss in final-frame prediction.
This amplification is visible in both representative regimes (Figure 4c). In the anisotropic branching sequence, final-frame IoU increases from 94.21% to 96.63%; in the near-concentric sequence, from 93.30% to 96.38%. This establishes a causal amplification mechanism: sub-percent-level segmentation differences are sufficient to induce multi-point degradation in long-horizon prediction. The mechanism is intrinsic to autoregressive boundary evolution, in which small geometric biases are recursively fed back as state and thereby converted into cumulative drift.
2.6 Temporal modeling of front dynamics
Temporal prediction must preserve two coupled properties: where the boundary lands (geometric fidelity) and how it gets there (dynamical consistency). The evaluation therefore spans eight complementary metrics (Figure 5b), with a deliberate hierarchy for interpretation: boundary-sensitive distances (HD, HD95, ASSD) and overlap (mIoU) report localization; propagation RMSE and TCI test trajectory-level stability; angular deviations ($|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$) diagnose whether anisotropic branching is preserved beyond coarse extent.
Figure 5: Morpher enables geometry-consistent long-horizon forecasting of swarming colony morphology. a Morpher architecture. Observed masks are encoded into a compact morphological latent sequence, and future evolution is predicted autoregressively with decoded states recursively conditioning subsequent steps. The Morphon module retrieves past states via cross-attention and integrates them through a learnable gate, while a multi-scale decoder preserves boundary detail and front geometry. b Evaluation of temporal architectures and inference strategies under an 80% observation / 20% prediction protocol. Autoregressive inference consistently outperforms parallel decoding, and Morphon improves boundary fidelity and temporal consistency across RNN, LSTM, GRU, and Transformer models. c Robustness to observation length. With Morphon enabled, prediction accuracy and stability improve as the observation ratio increases from 50% to 90%, indicating effective use of temporal context without overfitting.
We analyzed temporal architectures (RNN, GRU, LSTM, and Transformer) under parallel and autoregressive decoding, with or without Morphon (Figure 5b). Autoregressive decoding consistently outperforms parallel prediction. For Transformer with Morphon, performance improves from 94.80% mIoU and 4.63 px ASSD in parallel mode to 95.42% and 3.93 px under autoregressive decoding. GRU and LSTM show the same directional improvement. Stepwise state updates preserve local geometric continuity, whereas one-shot prediction smooths the front.
The backbones nonetheless separate along different axes of performance. The plain RNN is weakest: without Morphon in parallel mode, it reaches 93.23% mIoU, 6.02 px ASSD, and $|\Delta\mathrm{NAS}|=19.54\%$, producing overly round and contracted fronts. GRU is substantially stronger; even without Morphon in parallel mode, it reduces ASSD to 5.17 px and $|\Delta\mathrm{NAS}|$ to 15.31%, and in its best configuration it further lowers ASSD to 4.06 px and achieves the lowest propagation RMSE, 2.06 px per frame. LSTM is more stable than RNN and keeps RMSE between 2.26 and 2.84 px per frame, but reacts less strongly than GRU to fine front undulations.
Critically, the model that minimizes propagation error is not the one that best preserves morphology. GRU achieves the lowest RMSE (2.06 px per frame). Under the same setting, it remains worse than Transformer in boundary fidelity and anisotropic structure preservation. In the autoregressive Morphon setting, Transformer yields the highest mIoU (95.42%), the lowest ASSD (3.93 px), and the smallest anisotropy deviation ($|\Delta\mathrm{NAS}|=13.13\%$), while keeping both propagation RMSE and $|\Delta\mathrm{H}_{2}|$ low. Accurate front advance and faithful morphology reconstruction are therefore not identical objectives. The former is largely a gated-memory problem, whereas the latter requires long-range retrieval of branch history and spatial configuration.
Morphon improves every backbone, and its effect is diagnostic rather than cosmetic: it specifically targets the non-Markovian part of boundary evolution, where similar instantaneous curvatures can lead to different futures depending on earlier branch history. In parallel RNN, Morphon raises mIoU from 93.23% to 94.22% and reduces ASSD from 6.02 to 4.85 px. In autoregressive GRU, it reduces HD95 from 12.03 to 10.14 px. In Transformer, it lowers ASSD from 5.26 to 3.93 px and decreases $|\Delta\mathrm{NAS}|$ from 16.83% to 13.13%. Notably, $|\Delta\mathrm{H}_{2}|$ remains below 2.0% across models, suggesting that low-order harmonic structure is largely preserved even by weaker predictors. The main separation therefore lies not in coarse global shape, but in the preservation of finer anisotropic organization along the advancing front.
2.7 Scaling with observation length
The observation ratio was varied from 50% to 90% for each backbone in its best configuration, namely autoregressive decoding with Morphon (Figure 5c). We begin at 50% observation because the early phase is morphologically convergent across colonies (Figure 6d–f), and prediction only becomes well-posed once colony-specific geometry begins to diverge. Geometry-based metrics improve smoothly with longer history. For the Transformer, mIoU rises from 88.22% at 50% observation to 96.79% at 90%, while ASSD falls from 9.49 to 2.75 px. The relative ordering of the backbones remains unchanged across the full range, indicating that temporal architecture matters more than the exact observation–prediction split.
Front-propagation RMSE follows the same direction but saturates earlier. The largest gain appears between 50% and 60% observation; beyond that, values remain within a narrow 2.0–2.3 px per frame band. TCI is similarly stable, clustering around 63–65% from 50% to 80% observation, with the highest value observed for LSTM at 80% observation (65.32%). At 90%, the prediction window is too short for a stable TCI estimate.
Angular metrics expose a stricter regime and reveal a second constraint: more history does not guarantee better anisotropy preservation because the error is dominated by phase alignment of late-stage branch rearrangements. For the Transformer, $|\Delta\mathrm{NAS}|$ changes from 14.62% at 50% observation to 11.93% at 60%, rises to 15.57% at 70%, drops to 13.13% at 80%, and remains around 14% at 90%. $|\Delta\mathrm{H}_{2}|$ shows a different but equally late-stage-sensitive trend: for the Transformer it is 1.18–1.35% at 50–60% observation and approaches $\sim$2.2% at 90%. Additional history therefore yields diminishing returns for low-order shape statistics, while leaving phase-sensitive alignment of late front rearrangements as the dominant residual difficulty.
2.8 Generalization of trajectory-level dynamics
Out-of-sample stability was tested by leave-one-out cross-validation, evaluating each trajectory with models trained on all remaining trajectories (Figure 6). Performance approaches saturation as the training set expands, with only marginal gains once most trajectories are included (Figure 6a), indicating that the dominant modes of swarming variation relevant for forecasting are already well represented.
Figure 6: Generalization, accuracy, and dynamical consistency of swarming morphology forecasting under data-limited conditions. Results are evaluated using leave-one-out cross-validation across 81 independent colonies. a Effect of training set size on forecasting performance (80% observation / 20% prediction). Metrics converge with increasing data, indicating that the observed variability is sufficiently captured. b Static forecasting accuracy (80% / 20%) measured by region- and boundary-based metrics. c Dynamical consistency (80% / 20%) relating propagation error, anisotropy deviation, and normalized performance. d–f Time-aligned trajectories of effective radius, perimeter, and front propagation velocity under a 50% observation / 50% prediction protocol, with individual trajectories and mean trends. g Prediction-horizon–aligned trajectories (50% / 50%) showing forecast evolution within the prediction window. h Prediction-horizon–aligned trajectories (80% / 20%) showing consistent behavior across observation regimes.
Leave-one-out evaluation separates static accuracy from dynamical fidelity. Region- and boundary-based metrics remain concentrated in a high-performance regime (Figure 6b), whereas dynamical measures separate propagation error, anisotropy deviation, and harmonic distortion more clearly (Figure 6c). The model therefore generalizes well in mask overlap without collapsing the more sensitive dynamical dimensions into the same signal.
This separation is preserved at the trajectory level. Under a 50% observation / 50% prediction protocol, predicted effective radius and perimeter closely track the true time-aligned trajectories, and front propagation velocity reproduces the transition from rapid early expansion to later stabilization (Figure 6d–f). Agreement remains after alignment by prediction horizon (Figure 6g). Under an 80% observation / 20% prediction split, the same dynamical organization is retained (Figure 6h): longer observation tightens prediction over a shorter horizon without changing the overall structure of the trajectories. Across settings, Morpher preserves not only colony extent but also the temporal organization of expansion.
3 Discussion
Swarming is a striking example of emergent collective behavior [66, 25], yet its advancing front remains treated as a descriptive pattern rather than a predictive dynamical state. Here we recast Enterobacter sp. SM3 expansion as a predictive problem defined at the colony edge. This shift is biologically consequential: in inflamed intestinal environments, the advancing front is the interface through which swarming reshapes oxygen availability, niche structure, and downstream community organization (Figure 1[14]. The question is therefore not simply how a colony looks, but whether the future position and organization of its front can be inferred from its present state.
A central result is that this future is not cleanly indexed by nominal assay conditions. Trajectories remain broadly overlapping across temperature, humidity, and agar groupings (Figure 2a,b), and within- and across-condition distances are only weakly separated (Figure 2c,d). Prediction therefore depends on colony-specific morphology rather than on coarse labels. This helps explain why appearance-based extrapolation fails: the relevant information is carried by front geometry, protrusions, and temporal continuity, not by condition identity or framewise visual similarity [65].
That observation makes measurement decisive. TexPol–Net matters not simply because it improves segmentation, but because it defines the state on which dynamics are learned. Once the colony edge is treated as the evolving variable, small boundary errors are no longer cosmetic (Figure 3c–e). A sub-point difference in segmentation quality between TexPol–Net and YOLOv12 (92.48% vs. 91.81% mAP50:95) expands into a 2.4–3.1 IoU loss after autoregressive rollout (Figure 3c and Figure 4c). Stable forecasting therefore requires boundary fidelity at the measurement stage.
Within that state space, Morpher behaves, as schematized in Figure 5a, as a front-dynamics model rather than an image predictor. Its advantage over generic video architectures lies not only in overlap, but in preserving front localization and anisotropic branching over long horizons (Figure 4a,b). The temporal ablations sharpen this point. Autoregressive decoding consistently outperforms parallel prediction, indicating that swarming is better modeled as incremental state propagation than as one-shot future synthesis (Figure 5b). More importantly, the model that minimizes propagation error is not the one that best preserves morphology: GRU achieves the lowest RMSE, whereas Transformer with Morphon best maintains boundary fidelity and anisotropic structure (Figure 5b). The harder part of swarming prediction is therefore not coarse advance alone, but retention of branch history and late-stage front organization [67].
The observation-ratio and leave-one-out analyses place that difficulty more precisely. Global geometric metrics improve and then saturate with longer observation, whereas angular descriptors remain sensitive to late rearrangements of the front (Figure 5c). At the same time, leave-one-out evaluation shows that the model preserves not only mask overlap but also the temporal organization of effective radius, perimeter, and propagation velocity across held-out trajectories (Figure 6b,d–h). Together, these results indicate that the dominant residual challenge lies in anisotropic branching and phase-sensitive reorganization, not in recovery of coarse colony extent.
Several limitations define the next steps. While leave-one-out evaluation already indicates that the model captures transferable dynamics under limited data, SwarmEvo is currently restricted to a single strain, Enterobacter sp. SM3, with broader taxonomic and substrate diversity remaining as a natural extension to assess the generality of these dynamics. The present representation is quasi-two-dimensional, so height and density enter only indirectly. Morpher also operates without explicit biophysical constraints; incorporating mechanistic priors may further improve extrapolation in sparsely sampled regimes. Finally, forecasting remains open-loop. Closed-loop perturbation experiments guided by model predictions would provide the most direct test of whether predictive morphology can be translated into controllable intervention at the swarming front.
More broadly, this framework suggests a route for living systems in which shape carries the essential dynamics [19]. By separating measurement, state construction, and temporal evolution, it becomes possible to ask which aspects of collective behavior are predictable, which remain unstable, and which are most sensitive to perturbation. For swarming, this turns morphology from a descriptive readout into a dynamical variable that can be compared, forecasted, and ultimately acted upon.
4 Conclusions
Boundary-resolved morphology provides a workable state space for forecasting Enterobacter sp. SM3 swarming. SwarmEvo supplies the time-lapse trajectories, TexPol–Net recovers fronts with sufficient fidelity to define the predictive state, and Morpher advances those states through autoregressive rollout with structural memory. Under the 80% observation / 20% prediction protocol, Morpher achieves the best overall performance (mIoU 95.42%, HD95 10.61 px, ASSD 3.93 px), outperforming the strongest baseline by 5.4% in overlap and reducing HD95 and ASSD by 42.0% and 55.7%, respectively. Just as importantly, the study shows why forecasting is feasible only in a geometry-first formulation: once morphology is measured accurately, swarming expansion can be treated as a predictable dynamical process rather than an appearance-driven one. This establishes a basis for quantitative comparison of swarming trajectories across environments and for future predictive control of advancing microbial fronts.
5 Methods
5.1 Swarming assay, imaging, and dataset construction
The experimental workflow followed Figure 1. A single colony of Enterobacter sp. SM3 was transferred from an LB agar plate into LB broth (10 g/L tryptone, 5 g/L yeast extract, and 5 g/L NaCl) and cultured overnight at $37\,^{\circ}\mathrm{C}$ with shaking at 200 rpm. A $5\text{--}8\,\mu\mathrm{L}$ aliquot was inoculated at the center of a freshly prepared swarming plate (LB with 0.5%, 0.6%, or 0.7% agar; plate thickness 3–4 mm), incubated at $30\,^{\circ}\mathrm{C}$ and $\sim 90\%$ relative humidity for 4–6 h to activate swarming, and then transferred to a time-lapse imaging chamber maintained at $27\,^{\circ}\mathrm{C}$, $30\,^{\circ}\mathrm{C}$, or $33\,^{\circ}\mathrm{C}$ and $86\%$, $90\%$, or $94\%$ relative humidity. These temperature, humidity, and agar conditions span the permissive regime of SM3 swarming [47, 11, 27] and were varied across discrete levels to probe model generalization within physiologically relevant growth states. Humidity was kept below the condensation threshold to avoid droplet formation on the agar surface. Experimental conditions were varied to generate distinct expansion regimes.
Images were acquired with a vertically mounted high-resolution digital camera under uniform LED illumination. Frames were recorded every minute until the colony reached the plate boundary or no further measurable expansion was observed. The resulting Swarming Morphogenesis Evolution (SwarmEvo) dataset comprised 1,971 annotated images for segmentation and 276 long time series for temporal modeling. All recordings were stored at native resolution ($1250\times 1250$ px) with timestamps.
For downstream analysis, each sequence was converted into boundary-resolved colony masks using TexPol–Net. These masks defined the state space for forecasting. Training and validation partitions were split at the sequence level, with no colony contributing trajectories to both partitions. Exact dataset composition and augmentation procedures are provided in the Supplementary Information.
5.2 TexPol–Net for boundary-resolved state construction
TexPol–Net was designed to recover colony fronts under two coupled conditions: diffuse, low-contrast boundaries at the local scale and near-concentric radial organization at the colony scale. The network therefore combines two complementary modules. Texture–Edge Attention (TEA) preserves fine boundary texture through local depthwise filtering, multi-scale dilated texture encoding, and an edge-sensitive high-pass prior. Polar–Context Attention (PCA) embeds a geometry-aligned prior by combining local features, large-kernel Cartesian context, and a polar branch operating in $(\rho,\theta)$ coordinates. Full module formulations and implementation details are provided in the Supplementary Information.
As shown in Figure 3a, the full architecture follows a prototype-based one-stage instance segmentation design [52, 5]. A five-stage hierarchical convolutional backbone progressively downsamples the input while interleaving TEA and PCA blocks, allowing fine boundary cues and geometry-aligned context to be encoded jointly. The resulting multi-scale features are fused by a PANet-style bidirectional neck [36, 38], in which PCA is retained during top-down and bottom-up aggregation to preserve polar consistency across resolutions. Prediction is performed at feature levels $P3$, $P4$, and $P5$ through dense heads for class scores, bounding boxes, and instance-specific mask coefficients. In parallel, a lightweight Protonet produces $k=32$ shared prototypes, which are linearly combined with the predicted coefficients and then cropped and thresholded to produce final colony masks.
Training uses a composite segmentation objective,
$\mathcal{L}_{\mathrm{seg}}=\lambda_{b}\mathcal{L}_{b}+\lambda_{c}\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{m}\mathcal{L}_{m}.$
Here $\mathcal{L}_{b}$, $\mathcal{L}_{c}$, $\mathcal{L}_{d}$, and $\mathcal{L}_{m}$ denote box, classification, distribution focal, and mask losses, respectively. All weights were fixed across experiments.
5.3 Morpher for morphology forecasting
Morpher forecasts swarming expansion in mask space rather than image space. This formulation follows the biology of the problem: what propagates is the front, not visual appearance. Each input mask sequence is encoded by a shared multi-scale spatial encoder into framewise latent descriptors $z_{t}\in\mathbb{R}^{256}$, together with intermediate feature maps used later for decoding. Sinusoidal temporal encodings are added before temporal modeling.
As summarized in Figure 5a, Morpher consists of four coupled components: a multi-scale spatial encoder, a temporal sequence model, a Morphon memory block, and a multi-scale decoder. Observed masks are first compressed into a compact latent sequence, while encoder-side intermediate features are retained for later reconstruction. Future evolution is then predicted in latent space. In the autoregressive setting, each decoded prediction is re-encoded and fed back as the next input, so that the forecast proceeds through stepwise updates rather than one-shot generation. Morphon operates on the observed latent history by cross-attention with a learnable query derived from the aggregated observation state, and injects the retrieved structural memory through a learnable gate $\alpha\in(0,1)$. A multi-scale decoder finally reconstructs the predicted mask while reinjecting encoder features to preserve peripheral protrusions and fine curvature.
Forecasting is performed autoregressively unless otherwise stated. After observing $T_{\mathrm{obs}}$ frames, the model predicts the next latent state, decodes it into a mask, re-encodes that prediction, and feeds it back for subsequent steps. This stepwise rollout couples future predictions to the geometry produced at the previous step and stabilizes boundary continuity over long horizons. Parallel prediction is used only in matched ablations.
The temporal module is instantiated as one of four matched-capacity sequence models: vanilla RNN [16], GRU [13], LSTM [22], or Transformer encoder [59]. All variants share the same encoder–decoder backbone, the same latent dimensionality, and the same past-to-state formulation. The recurrent variants use three recurrent layers; the Transformer uses three encoder blocks.
Prediction is supervised at each forecast step by a temporally averaged objective,
$\mathcal{L}_{\mathrm{pred}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}^{(t)},\quad\mathcal{L}^{(t)}=\lambda_{f}\mathcal{L}_{f}+\lambda_{i}\mathcal{L}_{i}+\lambda_{b}\mathcal{L}_{b}.$
Here $\mathcal{L}_{f}$, $\mathcal{L}_{i}$, and $\mathcal{L}_{b}$ denote focal, soft IoU, and boundary losses, respectively. All weights were fixed across experiments. Implementation details, sequence construction, and baseline-matched comparisons are provided in the Supplementary Information.
5.4 Training and implementation
For TexPol–Net, all images were resized to $640\times 640$. Training used the Ultralytics YOLO framework with SGD ($6\times 10^{-3}$ initial learning rate, three-epoch linear warm-up, decay to 1% of the initial rate, momentum =0.937=0.937, weight decay $5\times 10^{-4}$), batch size =16=16, mixed-precision training, and early stopping once validation performance saturated.
For Morpher, binary masks generated by a single segmentation model were uniformly subsampled with a fixed stride and partitioned into observation and prediction segments under ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$. All masks were resized to $640\times 640$. Training used AdamW ($5\times 10^{-5}$ initial learning rate, weight decay $10^{-4}$), batch size =2=2, 300 epochs, 10% warm-up followed by cosine annealing, gradient clipping at 1.0, and mixed precision. Model selection was based on the highest validation mIoU. No early stopping was applied. All experiments used fixed random seeds and deterministic backend settings; TensorFloat-32 acceleration was enabled where available.
All baseline models were retrained under matched preprocessing and comparable train/validation splits. Model-specific implementation details, including optimizer schedules and architecture-specific settings for YOLOv11/12, SAM/SAM2, MAU, MIM, PredRNN, PredRNNv2, SimVP+TAU, and SimVPv2+gSTA, are provided in the Supplementary Information.
5.5 Evaluation
Segmentation was evaluated by mAP50:95, image-wise IoU, and Dice coefficient. Forecasting was evaluated at two levels. Spatial fidelity was assessed by mIoU, HD, HD95, and ASSD, which quantify overlap and boundary localization. Dynamical fidelity was assessed by radial-velocity RMSE and the Temporal Consistency Index (TCI), which measure front propagation and temporal fluctuation preservation. Angular organization was quantified by $|\Delta\mathrm{NAS}|$ and $|\Delta\mathrm{H}_{2}|$, which measure deviations in anisotropic growth and second-harmonic angular structure.
In the main text, mAP50:95 serves as the primary segmentation metric, while HD95 and ASSD serve as the primary forecasting metrics because region overlap can remain high even when the contour drifts. Formal definitions of all metrics are provided in the Supplementary Information.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC, Grant No. 12202275) and the Shanghai Jiao Tong University Explore X Fund.
Conflict of Interest
The authors declare no conflict of interest.
Data availability
All datasets used in this study are publicly available. The Swarming Morphogenesis Evolution (SwarmEvo) dataset, including both the segmentation and temporal prediction subsets, is hosted at https://huggingface.co/datasets/ShengyouDuan/SwarmEvo. The dataset provides polygon-based segmentation annotations and time-lapse sequences for bacterial swarming experiments, and is released to support reproducibility and further research in morphology-aware modeling.
Code availability
The complete implementation of the proposed framework, including TexPol–Net for morphology-aware segmentation and Morpher for autoregressive temporal forecasting, is publicly available at https://github.com/ShengyouDuan/From_shape_to_fate__making_bacterial_swarming_expansion_predictable. The repository contains all training and evaluation code required to reproduce the results reported in this work.
Supplementary Information
S1 Texture–Edge Attention and Polar–Context Attention modules
Swarming colony images exhibit complex morphological organization characterized by uncertain boundaries, irregular shapes, and radially propagating texture patterns. These inherent properties pose significant challenges for CNNs, whose reliance on local receptive fields restricts their capacity to capture long-range spatial dependencies and global geometric structures, particularly the near-concentric-ring radial expansion typical of swarming growth. To overcome these limitations and enhance both fine-grained texture extraction and high-level semantic representation, we developed two specialized attention modules, the Texture–Edge Attention (TEA) and Polar–Context Attention (PCA), as illustrated in Figure 7.
The TEA module, as shown in Figure 7a, is designed to address blurred boundaries and multi-scale, high-frequency texture variability. It combines three cooperative paths: a local branch that preserves intra-channel spatial details, a multi-dilated path that ensures scale-robust texture encoding, and an edge-sensitive path initialized with a discrete Laplacian kernel to enhance boundary awareness. Channel-wise and spatial gating mechanisms further refine the fused representation by emphasizing informative structures while maintaining computational efficiency.
Let the input be $\mathbf{X}\!\in\!\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ and the output be $\mathbf{Y}\!\in\!\mathbb{R}^{B\times C_{\mathrm{out}}\times H\times W}$, where $B$ is the batch size, $C_{\mathrm{in}}$ and $C_{\mathrm{out}}$ denote the number of input and output channels, and $H$ and $W$ represent spatial dimensions. To balance representation capacity and computational cost, an intermediate channel width is introduced as
$C_{h}=C_{\mathrm{out}}\cdot e,$ (1)
where $e\in(0,1]$ is the expansion ratio controlling internal dimensionality.
Local features are first extracted using a depthwise $3\times 3$ convolution to capture intra-channel spatial structures, followed by a $1\times 1$ pointwise projection to $C_{h}$ channels to ensure dimensional consistency. The normalized and activated local features are
$\mathbf{F}_{\mathrm{loc}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\left(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}(\mathbf{X})\right)\right)\right),$ (2)
where $\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}$ denotes depthwise convolution, $\mathrm{GN}$ represents group normalization, and $\phi$ is the SiLU activation.
A squeeze-and-excitation (SE) gate $\mathbf{f}\!\in\!\mathbb{R}^{B\times C_{h}\times 1\times 1}$ is computed via global average pooling (GAP) followed by two $1\times 1$ convolutions with nonlinearity and sigmoid activation:
$\mathbf{f}=\sigma\!\Big(\mathrm{Conv}_{1\times 1}\big(\phi(\mathrm{Conv}_{1\times 1}(\mathrm{GAP}(\mathbf{X})))\big)\Big).$ (3)
Channel gating is applied element-wise:
$\widetilde{\mathbf{F}}_{\mathrm{loc}}(c,h,w)=\mathbf{F}_{\mathrm{loc}}(c,h,w)\odot\mathbf{f}(c).$ (4)
To model textures across multiple scales, a multi-dilated branch applies depthwise convolutions with dilation factors $d_{k}\!\in\!\mathcal{D}=\{d_{1},d_{2},\ldots,d_{K}\}$:
$\mathbf{F}^{(d_{k})}_{\mathrm{tex}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\Big(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3,d_{k}}(\mathbf{X})\Big)\right)\right),$ (5)
Figure 7: Texture–Edge Attention (TEA) and Polar–Context Attention (PCA) modules. a, The TEA block enhances fine-scale texture fidelity and boundary sharpness through three cooperative branches: a local depthwise path for intra-channel spatial preservation, multi-dilated convolutions for scale-robust texture encoding, and an edge-sensitive Laplacian path that injects a high-pass prior. Channel and spatial gating further refine feature fusion, producing an edge-aware, redundancy-suppressed representation. b, The PCA block embeds a polar-aware geometric prior aligned with the radial growth of swarming colonies. Input features are first compressed and then processed by a local branch, a large-kernel Cartesian branch, and a polar-warped branch operating in $(\rho,\theta)$ coordinates. Depthwise dilated filters extract context along radial and angular axes, and subsequent channel- and spatial-attention gates yield a geometry-aligned output.
The concatenation of all paths yields
$\mathbf{F}_{\mathrm{tex}}=\mathrm{Concat}\big(\mathbf{F}^{(d_{1})}_{\mathrm{tex}},\mathbf{F}^{(d_{2})}_{\mathrm{tex}},\ldots,\mathbf{F}^{(d_{K})}_{\mathrm{tex}}\big)$, (6)
where K=3K{=}3 captures short-, medium-, and long-range textures.
An edge-aware branch initialized by the Laplacian kernel enhances boundary sensitivity:
$\mathbf{F}_{\mathrm{edge}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\Big(\mathrm{Conv}^{\mathrm{dw,Lap}}_{3\times 3}(\mathbf{X})\Big)\right)\right),$ (7)
with initialization
$\mathbf{K}_{\mathrm{Lap}}=\begin{bmatrix}0&-1&0\\[-2.0pt] -1&4&-1\\[-2.0pt] 0&-1&0\end{bmatrix}$. (8)
The texture and edge features are concatenated and reweighted by an SE gate:
$\mathbf{F}_{\mathrm{tex+edge}}=\mathrm{Concat}(\mathbf{F}_{\mathrm{tex}},\,\mathbf{F}_{\mathrm{edge}}),$ (9)
$\widetilde{\mathbf{F}}_{\mathrm{tex+edge}}=\mathbf{F}_{\mathrm{tex+edge}}\odot\sigma\!\Big(\mathrm{Conv}_{1\times 1}\big(\phi(\mathrm{Conv}_{1\times 1}(\mathrm{GAP}(\mathbf{F}_{\mathrm{tex+edge}})))\big)\Big).$ (10)
The SE-weighted outputs are combined with local features:
$\mathbf{F}_{\mathrm{fuse}}=\mathrm{Concat}(\widetilde{\mathbf{F}}_{\mathrm{loc}},\,\widetilde{\mathbf{F}}_{\mathrm{tex+edge}}).$ (11)
A spatial attention gate $\mathbf{g}\!\in\!\mathbb{R}^{B\times 1\times H\times W}$ emphasizes salient regions:
$\mathbf{g}=\sigma\!\Big(\mathrm{Conv}_{1\times 1}\big(\phi(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}(\mathbf{F}_{\mathrm{fuse}}))\big)\Big),$ (12)
and is applied element-wise:
$\bar{\mathbf{F}}_{\mathrm{fuse}}(c,h,w)=\mathbf{F}_{\mathrm{fuse}}(c,h,w)\odot\mathbf{g}(h,w).$ (13)
The fused representation is projected to the output dimension:
$\mathbf{Y}^{\prime}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}(\bar{\mathbf{F}}_{\mathrm{fuse}})\right)\right),$ (14)
with a conditional residual connection:
$\mathbf{Y}=\begin{cases}\mathbf{X}+\boldsymbol{\gamma}\odot\mathbf{Y}^{\prime},&C_{\mathrm{in}}=C_{\mathrm{out}},\\[3.0pt] \mathbf{Y}^{\prime},&C_{\mathrm{in}}\neq C_{\mathrm{out}},\end{cases}$ (15)
where $\boldsymbol{\gamma}\!\in\!\mathbb{R}^{C_{\mathrm{out}}}$ is a learnable scaling factor that ensures stability when the dimensions are the same.
While TEA focuses on boundary and texture fidelity, the PCA module illustrated in Figure 7b captures long-range dependencies and radial geometric organization inherent to swarming colonies. Conventional convolutional operators struggle with near-concentric-ring propagation whereas PCA embeds a polar-aware representation aligned with colony growth.
The module comprises three paths: a local branch for spatial detail, a large-kernel branch for contextual encoding, and a polar branch that transforms features into polar coordinates for radial modeling. The local branch follows the same $3\times 3$ depthwise-pointwise pattern described in Eq. 2, operating on $\mathbf{X}^{\prime}$:
$\mathbf{F}_{\mathrm{local}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\big(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3}(\mathbf{X}^{\prime})\big)\right)\right).$
Input $\mathbf{X}\!\in\!\mathbb{R}^{B\times C_{\mathrm{in}}\times H\times W}$ is first compressed by a $1\times 1$ convolution to $C_{h}$ channels, normalized, and activated to yield $\mathbf{X}^{\prime}$. The large-context branch employs a depthwise separable $7\times 7$ convolution:
$\mathbf{F}_{\mathrm{large}}=\phi\!\left(\mathrm{GN}\!\left(\mathrm{Conv}_{1\times 1}\!\Big(\mathrm{Conv}^{\mathrm{dw}}_{7\times 7}(\mathbf{X}^{\prime})\Big)\right)\right).$ (16)
Spatial indices $(h,w)$ are mapped to polar coordinates:
$\theta_{w}$ $=\tfrac{2\pi w}{W},$ $w$ $\in\{0,\ldots,W{-}1\}$ (17)
$\rho_{h}$ $=\tfrac{h}{H-1},$ $h$ $\in\{0,\ldots,H{-}1\}$ (18)
and then transformed to normalized Cartesian coordinates:
$u(h,w)$ $=\rho_{h}\cos\theta_{w}$ (19)
$v(h,w)$ $=\rho_{h}\sin\theta_{w}$ (20)
Bilinear interpolation provides the polar-warped feature map:
$\mathbf{X}^{\mathrm{pol}}_{b,c,h,w}=\mathcal{I}\!\big(\mathbf{X}^{\prime}_{b,c,:,:};\,u(h,w),v(h,w)\big),\quad b\!\in\!\{0,\ldots,B{-}1\},\;c\!\in\!\{0,\ldots,C_{h}{-}1\},$ (21)
where $\mathcal{I}$ denotes the bilinear interpolation operator sampling $\mathbf{X}^{\prime}$ at polar coordinates $(u,v)$.
A depthwise $3\times 3$ convolution with dilation $d_{\mathrm{pol}}{=}4$ extracts polar-domain context:
$\mathbf{F}_{\mathrm{polar}}=\phi\!\left(\mathrm{GN}\!\Big(\mathrm{Conv}_{1\times 1}\big(\mathrm{Conv}^{\mathrm{dw}}_{3\times 3,\,4}(\mathbf{X}^{\mathrm{pol}})\big)\Big)\right).$ (22)
Finally, outputs from the three branches are concatenated:
$\mathbf{F}_{\mathrm{cat}}=\mathrm{Concat}(\mathbf{F}_{\mathrm{local}},\,\mathbf{F}_{\mathrm{large}},\,\mathbf{F}_{\mathrm{polar}}),$ (23)
followed by channel SE and spatial attention (Eq. 313) to produce $\widetilde{\bar{\mathbf{F}}}_{\mathrm{cat}}$. Projection and conditional residual (Eq. 1415) yield the final PCA output. This design preserves fine structural details, integrates global context, and explicitly embeds radial priors, enabling robust modeling of colony expansion dynamics.
S2 Swarming Morphogenesis Evolution dataset
The Swarming Morphogenesis Evolution (SwarmEvo) dataset consists of high-resolution time-lapse recordings of Enterobacter sp. SM3 acquired at a fixed spatial resolution of $1250\times 1250$ px. Swarming colonies were initiated by inoculating a $5\text{--}8\,\mu\mathrm{L}$ aliquot from an overnight culture (grown at $37\,^{\circ}\mathrm{C}$, 200 rpm in LB broth) onto the center of freshly prepared LB agar plates with concentrations of 0.5%, 0.6%, or 0.7% (plate thickness 3–4 mm). Plates were first incubated at $30\,^{\circ}\mathrm{C}$ and approximately 90% relative humidity for 4–6 h to activate swarming, and subsequently transferred to a time-lapse imaging chamber with controlled environmental conditions. During imaging, temperature was set to $27\,^{\circ}\mathrm{C}$, $30\,^{\circ}\mathrm{C}$, or $33\,^{\circ}\mathrm{C}$, and relative humidity to 86%, 90%, or 94%, spanning the permissive regime of SM3 swarming. Humidity was maintained below the condensation threshold to prevent droplet formation on the agar surface. These controlled variations in agar concentration, temperature, and humidity produced distinct expansion regimes and were used to probe model generalization across physiologically relevant growth states.
After augmentation, the dataset comprises 1,971 annotated samples used for training and evaluating segmentation models, as well as 276 long time series derived from continuous recordings sampled at 1-min intervals, which serve as the basis for temporal modeling and multi-scale temporal downsampling. Data were collected across multiple agar plates and independent imaging sessions, introducing natural variability in growth dynamics and colony morphology. Segmentation masks used for model training and evaluation were obtained through a dedicated segmentation pipeline and subsequently curated to ensure consistent boundary delineation, while temporal sequences were generated by propagating these masks across time to support forecasting tasks.
Segmentation-level augmentation. For training the segmentation model, augmentations were applied independently to each image–annotation pair. Photometric perturbations included linear intensity rescaling with offset, gamma correction, additive Gaussian noise, and sparse impulse-like pixel corruption. Geometric transformations were sampled per image and applied consistently to the image and its polygon annotations, including random in-plane rotation, isotropic scaling, translation constrained by the instance extent, and random horizontal or vertical flipping. A random cutout was further used to simulate partial occlusion; polygon annotations were updated by geometrically clipping the visible region and retaining valid connected components. After each transformation, polygon validity was enforced by automatic closure and self-intersection repair, and invalid or degenerate shapes were discarded.
Sequence-level augmentation. Data augmentation was applied at the sequence level and restricted to spatial transformations that preserve the underlying growth dynamics. For each sequence, a single set of affine transformation parameters was sampled and applied identically to all frames to maintain temporal coherence. The augmentation pipeline was limited to in-plane rotation, translation, and random horizontal and vertical flipping. No augmentation was applied selectively to specific temporal segments or across time. Temporal resolution was defined solely by fixed-stride subsampling, without stochastic temporal perturbations.
S3 Implementation of TexPol–Net
Framework and experimental setup. All models were trained using the Ultralytics YOLO framework, with the task configured for image segmentation. All experiments were conducted under identical training configurations and independently repeated to ensure fair comparability.
Training and optimization protocol. For training, the maximum number of epochs was set to 300, and an early stopping strategy was employed to mitigate overfitting once validation performance saturated. The batch size was set to 16, and all input images were resized to a fixed resolution of $640\times 640$, balancing computational efficiency and memory usage. Automatic mixed-precision training was enabled throughout to improve training throughput and reduce memory consumption while maintaining numerical stability. Optimization used stochastic gradient descent with an initial learning rate of $6\times 10^{-3}$, combined with a linear warm-up schedule over the first three epochs. After warm-up, the learning rate was gradually decayed to $1\%$ of its initial value. Weight decay and momentum were set to $5\times 10^{-4}$ and $0.937$, respectively. All experiments used a fixed random seed and deterministic training settings to ensure reproducibility.
S4 Implementation of Morpher
Sequence construction and data preprocessing. Morpher operated exclusively on binary colony masks generated by a pretrained segmentation model and did not directly access raw image intensities. In the primary pipeline, TexPol–Net was used to generate these masks, while alternative segmentation models were used in comparative experiments. Each training sample therefore consisted of a temporally ordered sequence of segmentation masks representing colony occupancy and front geometry. Sequences were constructed by uniformly subsampling frames from the full temporal series using a fixed stride, yielding equal temporal spacing between adjacent frames, and the resulting sequence length $T$ was defined by this fixed-stride subsampling rule. Each sequence was partitioned into an observation segment and a prediction segment using observation–prediction ratios of $0.5/0.5$, $0.6/0.4$, $0.7/0.3$, $0.8/0.2$, and $0.9/0.1$. Within each experiment, all masks were generated using the same segmentation model to ensure a consistent morphological representation across both segmentation and forecasting stages. All masks were resized to a spatial resolution of $640\times 640$, which was used consistently across all experiments. The dataset was split into training and validation partitions at the growth-sequence level, with no colony contributing sequences to both partitions, and all quantitative results were reported on the validation split.
Training and optimization protocol. All variants were optimized with AdamW using an initial learning rate of $5\times 10^{-5}$ and weight decay of $10^{-4}$. Training was conducted for 300 epochs with a batch size of 2. A linear warm-up was applied over the first $10\%$ of optimization steps, followed by cosine annealing. The global gradient norm was clipped to 1.0. Mixed-precision training was enabled throughout via automatic mixed precision with gradient scaling to improve computational throughput while maintaining numerical stability. Validation was performed at every epoch, and model selection was based on the checkpoint achieving the highest validation mIoU. No early stopping was applied. All experiments were conducted with fixed random seeds and deterministic backend settings to ensure reproducibility. TensorFloat-32 acceleration was enabled for matrix multiplications on supported hardware, while cuDNN was configured in deterministic mode.
S5 Running of existing methods
Adjustable training settings were kept aligned across methods whenever applicable. All images were resized to $640\times 640$. Models trained with epoch-based schedules were optimized for 300 epochs, while models trained with iteration-based schedules explicitly report the corresponding iteration counts. For all video prediction models, the batch size was fixed at 2.
YOLOv11 and YOLOv12. YOLOv11 and YOLOv12 were trained and evaluated using the Ultralytics YOLO framework with the task configured for image segmentation. Early stopping was enabled once validation performance saturated. A batch size of 16 was used. AMP was enabled throughout training. Optimization used SGD with an initial learning rate of $6\times 10^{-3}$ and a linear warm-up over the first three epochs; the learning rate was then decayed to $1\%$ of its initial value. Weight decay and momentum were set to $5\times 10^{-4}$ and $0.937$, respectively.
SAM and SAM2. SAM and SAM2 were fine-tuned for segmentation under a unified training protocol. For SAM, training used AdamW with an initial learning rate of $8\times 10^{-4}$ and weight decay $10^{-4}$. A warm-up phase of 250 optimization steps was applied at the beginning of training, followed by stepwise learning-rate decays at 60,000 and 86,666 iterations, each with a decay factor of $1/10$. The model was built on the SAM ViT-B backbone, and a selective freezing strategy was adopted: the image encoder and prompt encoder were frozen, while only the mask decoder was updated during training. For SAM2, training was formulated as a binary segmentation task with RGB images as input and binary masks as supervision. Images were converted to tensors and normalized to $[0,1]$, while masks were resized using nearest-neighbor interpolation and explicitly binarized. Optimization used Adam with an initial learning rate of $10^{-4}$, batch size 4, and binary cross-entropy loss with logits, without an additional learning-rate schedule.
MAU. MAU was run using the official implementation. The model employed four recurrent layers with hidden dimension 64, convolutional filters of size 5 with stride 1 and patch size 1, and no layer normalization. The spatiotemporal relation size was set to 2 and the temporal decay parameter to $\tau=5$. Scheduled sampling was enabled, with the sampling probability linearly decayed from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$. Training used Adam with a learning rate of $5\times 10^{-4}$ and a OneCycle learning-rate scheduler.
MIM. MIM was run using the official implementation built on the PredRNN framework. The model employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 4, and no layer normalization. Scheduled sampling followed the same linear decay strategy as above, while reverse scheduled sampling was disabled. Training used Adam with a learning rate of $10^{-4}$ and a OneCycle learning-rate scheduler; incomplete batches were dropped during training.
PredRNN and PredRNNv2. PredRNN-based models were run using the official implementations. Both models employed four recurrent layers with hidden dimensions of 128, convolutional filters of size 5 with stride 1 and patch size 2, and no layer normalization. Scheduled sampling was enabled with linear decay from 1.0 to 0 over 50,000 iterations at a rate of $2\times 10^{-5}$. Training used Adam with a learning rate of $10^{-3}$ and a OneCycle learning-rate scheduler. PredRNNv2 additionally enabled reverse scheduled sampling with transition steps at 25,000 and 50,000 iterations and an exponential coefficient of 5,000, and incorporated a decoupling loss with weight $\beta=0.01$.
SimVP and SimVPv2. SimVP-based baselines were run using the official implementation. The spatial encoder–decoder employed a channel width of 64 with four convolutional blocks ($N_{S}=4$), while temporal modeling used a hidden dimension of 256 with eight temporal blocks ($N_{T}=8$). SimVP used TAU units for temporal prediction, whereas SimVPv2 replaced TAU with gSTA modules. Training used Adam with a learning rate of $10^{-3}$ and a OneCycle learning-rate scheduler. Model selection followed the validation loss criterion defined in the configuration.
S6 Evaluation metrics for colony front segmentation
Segmentation performance was evaluated using three complementary metrics: mAP50:95, image-wise IoU, and Dice coefficient. mAP50:95 served as the primary segmentation metric because it summarizes performance across a range of localization tolerances. Following the COCO protocol,
$\mathrm{mAP}_{50:95}=\frac{1}{10}\sum_{\tau\in\{0.50,0.55,\dots,0.95\}}\left[\frac{1}{101}\sum_{i=0}^{100}P\!\left(\frac{i}{100};\tau\right)\right],$ (24)
where $P(r;\tau)$ denotes the interpolated precision at recall level $r$ under IoU threshold $\tau$. Dice was computed as
$\mathrm{Dice}=\frac{2\,|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|},$ (25)
where $\hat{Y}$ and $Y$ denote predicted and ground-truth masks.
S7 Evaluation metrics for morphological forecasting
Forecasting accuracy must be judged not only by per-frame agreement, but also by how faithfully the predicted colony advances and organizes its growth direction over time. Accordingly, evaluation considered both the spatial fidelity of the predicted masks and boundaries, and the temporal consistency of the advancing front, including its radial expansion speed and the coherence of its directional variation around the colony rim.
To quantify mask fidelity over time, let $\widehat{Y}_{t}$ and $Y_{t}$ denote the predicted and true masks at time $t$, and let $\partial\widehat{Y}_{t}$ and $\partial Y_{t}$ denote their corresponding boundaries. The mean Intersection over Union was defined as
$\mathrm{mIoU}=\frac{1}{T}\sum_{t=1}^{T}\frac{|\widehat{Y}_{t}\cap Y_{t}|}{|\widehat{Y}_{t}\cup Y_{t}|}.$ (26)
Boundary placement was evaluated using the symmetric Hausdorff distance,
$d_{\mathrm{H}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{\max_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|,\;\max_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\Big\},$ (27)
its 95th-percentile variant,
$d_{\mathrm{H}_{95}}=\frac{1}{T}\sum_{t=1}^{T}\max\Big\{P_{95}\!\big(\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|\big)_{y\in\partial Y_{t}},\;P_{95}\!\big(\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\big)_{\hat{y}\in\partial\widehat{Y}_{t}}\Big\},$ (28)
and the average symmetric surface distance,
$d_{\mathrm{ASSD}}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{|\partial\widehat{Y}_{t}|+|\partial Y_{t}|}\left(\sum_{y\in\partial Y_{t}}\min_{\hat{y}\in\partial\widehat{Y}_{t}}\|y-\hat{y}\|+\sum_{\hat{y}\in\partial\widehat{Y}_{t}}\min_{y\in\partial Y_{t}}\|\hat{y}-y\|\right).$ (29)
To evaluate whether the predicted colony reproduced the correct front propagation dynamics, radial expansion speed was measured along $K$ uniformly sampled angular directions $\{\theta_{k}\}_{k=1}^{K}$ from the colony centroid. In the experiments, $K=720$, corresponding to a sampling interval of $0.5^{\circ}$. Let $r(\theta_{k},t)$ and $\widehat{r}(\theta_{k},t)$ denote the ground-truth and predicted radial distances at time $t$. The corresponding expansion velocities were
$v(\theta_{k},t)=\frac{r(\theta_{k},t)-r(\theta_{k},t-\Delta t)}{\Delta t},\qquad\widehat{v}(\theta_{k},t)=\frac{\widehat{r}(\theta_{k},t)-\widehat{r}(\theta_{k},t-\Delta t)}{\Delta t}.$ (30)
The overall accuracy of front advancement was quantified by
$\mathrm{RMSE}=\frac{1}{T-1}\sum_{t=2}^{T}\sqrt{\frac{1}{K}\sum_{k=1}^{K}\big(\widehat{v}(\theta_{k},t)-v(\theta_{k},t)\big)^{2}}.$ (31)
Temporal fluctuation preservation was quantified by the Temporal Consistency Index (TCI) over sliding windows $w=1,\dots,W$ of fixed length $L=4$ radius frames, corresponding to three consecutive velocity steps. For each window and direction, $\sigma^{(w)}_{\widehat{v},k}$ and $\sigma^{(w)}_{v,k}$ denote the temporal standard deviations of predicted and ground-truth velocity traces. The directional consistency score was
$\mathrm{TCI}^{(w)}_{k}=1-\frac{\big|\sigma^{(w)}_{\widehat{v},k}-\sigma^{(w)}_{v,k}\big|}{\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}+\varepsilon},$ (32)
and was evaluated only when $\sigma^{(w)}_{\widehat{v},k}+\sigma^{(w)}_{v,k}>\tau_{0}$, where $\tau_{0}=10^{-6}$ filters directions with effectively no motion. The final index was
$\mathrm{TCI}=\frac{1}{W}\sum_{w=1}^{W}\frac{1}{|\mathcal{K}_{w}|}\sum_{k\in\mathcal{K}_{w}}\mathrm{TCI}^{(w)}_{k}$, (33)
where $\mathcal{K}_{w}$ denotes the set of valid directions in window $w$.
To evaluate the organization of growth across angles, the normalized angular spread (NAS) was computed from the angular standard deviation divided by the angular mean of the velocity field. We report the mean absolute deviation over time,
$|\Delta\mathrm{NAS}|=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{NAS}}_{t}-\mathrm{NAS}_{t}\big|.$ (34)
To further characterize directional patterning, we examined the angular Fourier spectrum of $v(\theta,t)$ and extracted the normalized second-harmonic power,
$\mathrm{H}_{2,t}\mathrm{H}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{v(\theta,t)\}[m]\big|^{2}},\qquad\widehat{\mathrm{H}}_{2,t}=\frac{\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[2]\big|^{2}}{\sum_{m}\big|\mathcal{F}_{\theta}\{\widehat{v}(\theta,t)\}[m]\big|^{2}}.$ (35)
The corresponding deviation was defined as
$|\Delta\mathrm{H}_{2}|=\frac{1}{T-1}\sum_{t=2}^{T}\big|\widehat{\mathrm{H}}_{2,t}-\mathrm{H}_{2,t}\big|.$ (36)
S8 Performance comparison with state-of-the-art video prediction models under an 80%–20% observation–prediction split
Table 1: Performance comparison with state-of-the-art video prediction models under an 80%–20% observation–prediction split. This table benchmarks Morpher against leading video prediction architectures, including MAU, MIM, PredRNN variants, and SimVP-based models. All methods are evaluated under identical input–output protocols for long-term forecasting of swarming colony expansion. Morpher achieves substantially higher region-level overlap (mIoU) and lower boundary error (HD95, ASSD), indicating improved accuracy in front propagation and boundary-level morphology.
Model mIoU (%) $\uparrow$ HD95 (px) $\downarrow$ ASSD (px) $\downarrow$
MAU 84.67 22.73 14.68
MIM 89.32 20.17 10.30
PredRNN 84.60 22.75 14.81
PredRNNv2 84.14 23.24 15.04
SimVP+TAU 86.87 23.19 12.47
SimVP+gSTA 90.52 18.28 8.87
Morpher (Ours) 95.42 10.61 3.93
S9 Performance of Morpher under an 80%–20% observation–prediction split across temporal modeling and inference paradigms
Table 2: Performance of Morpher under an 80%–20% observation–prediction split across temporal modeling and inference paradigms. Forecasting accuracy is evaluated across region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front-propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$). Higher mIoU and TCI indicate superior forecasting performance, whereas lower HD-based distances, RMSE, $\lvert\Delta\mathrm{NAS}\rvert$, and $\lvert\Delta\mathrm{H}_{2}\rvert$ reflect improved geometric and dynamical fidelity. This table provides a mechanistic comparison by isolating the effects of temporal modeling, inference strategy, and the Morphon memory mechanism.
Seq. Model Inference Paradigm Morphon mIoU (%) $\uparrow$ HD (px) $\downarrow$ HD95 (px) $\downarrow$ ASSD (px) $\downarrow$ RMSE (px/frame) $\downarrow$ TCI (%) $\uparrow$ $\lvert\Delta\mathrm{NAS}\rvert$ (%) $\downarrow$ $\lvert\Delta\mathrm{H}_{2}\rvert$ (%) $\downarrow$
RNN Parallel 93.23 17.68 12.85 6.02 3.36 55.34 19.54 1.96
LSTM Parallel 93.24 17.44 12.65 5.85 2.84 60.48 15.99 1.87
GRU Parallel 93.96 17.14 12.51 5.17 3.02 54.80 15.31 1.79
Transformer Parallel 93.94 17.24 12.64 5.23 2.95 56.94 16.83 1.94
RNN Autoregressive 93.55 17.53 12.51 5.26 2.20 65.63 15.02 1.90
LSTM Autoregressive 94.07 16.92 11.85 5.34 2.60 63.14 17.80 1.91
GRU Autoregressive 94.20 17.25 12.03 5.01 2.39 64.10 19.02 1.84
Transformer Autoregressive 94.16 16.56 11.95 5.26 2.66 64.92 13.51 1.74
RNN Parallel 94.22 16.67 11.63 4.85 2.85 61.57 17.57 1.84
LSTM Parallel 94.44 15.97 10.90 4.59 2.76 59.29 14.72 1.69
GRU Parallel 94.58 15.86 11.46 4.81 2.68 63.14 15.56 1.76
Transformer Parallel 94.80 15.79 11.22 4.63 2.55 62.71 16.30 1.89
RNN Autoregressive 94.94 15.46 10.67 4.34 2.19 63.32 16.87 1.86
LSTM Autoregressive 95.01 15.32 10.77 4.20 2.26 65.32 15.03 1.79
GRU Autoregressive 95.29 15.01 10.14 4.06 2.06 64.31 15.45 1.89
Transformer Autoregressive 95.42 15.26 10.61 3.93 2.12 64.26 13.13 1.85
S10 Performance of Morpher under a series of observation–prediction splits across sequence models
Table 3: Performance of Morpher under a series of observation–prediction splits across sequence models. Results are reported for 50%, 60%, 70%, 80%, and 90% observation levels to assess how forecasting stability changes as more of the past is revealed. Metrics include region-level overlap (mIoU), boundary accuracy (HD, HD95, ASSD), front-propagation dynamics (RMSE), temporal fluctuation consistency (TCI), and angular growth organization ($\lvert\Delta\mathrm{NAS}\rvert$, $\lvert\Delta\mathrm{H}_{2}\rvert$). Higher mIoU and TCI indicate superior forecasting performance, whereas lower HD-based distances, RMSE, $\lvert\Delta\mathrm{NAS}\rvert$, and $\lvert\Delta\mathrm{H}_{2}\rvert$ reflect improved geometric and dynamical fidelity.
Observation (%) Seq. Model mIoU (%) $\uparrow$ HD $\downarrow$ HD95 $\downarrow$ ASSD $\downarrow$ RMSE (px/frame) $\downarrow$ TCI (%) $\uparrow$ $\lvert\Delta\mathrm{NAS}\rvert$ (%) $\downarrow$ $\lvert\Delta\mathrm{H}_{2}\rvert$ (%) $\downarrow$
50 RNN 87.88 25.38 20.92 10.19 2.56 62.77 13.42 1.14
LSTM 88.18 25.15 20.37 9.43 2.20 62.97 14.28 1.21
GRU 87.99 25.89 21.35 10.02 2.42 62.71 14.37 1.15
Transformer 88.22 23.64 19.12 9.49 2.28 63.79 14.62 1.18
60 RNN 92.20 19.76 15.35 6.29 2.10 64.09 12.04 1.33
LSTM 91.76 20.26 15.38 6.71 2.22 64.34 14.46 1.38
GRU 92.24 19.48 15.30 6.33 2.03 64.13 12.98 1.37
Transformer 92.64 18.21 14.18 5.96 1.96 63.66 11.93 1.35
70 RNN 93.18 18.02 13.34 5.64 2.29 64.89 17.55 1.54
LSTM 93.40 18.58 13.57 5.41 2.18 64.54 14.64 1.47
GRU 93.37 18.57 13.78 5.59 2.25 64.49 16.83 1.52
Transformer 93.80 16.88 12.18 4.95 2.12 62.42 15.57 2.25
80 RNN 94.94 15.46 10.67 4.34 2.19 63.32 16.87 1.86
LSTM 95.01 15.32 10.77 4.20 2.26 65.32 15.03 1.79
GRU 95.29 15.01 10.14 4.06 2.06 64.31 15.45 1.89
Transformer 95.42 15.26 10.61 3.93 2.12 64.26 13.13 1.85
90 RNN 96.02 12.48 8.36 3.13 2.24 13.24 2.20
LSTM 96.31 12.45 8.19 2.95 2.08 13.59 2.33
GRU 96.42 12.09 8.38 3.07 2.09 13.86 2.24
Transformer 96.79 11.20 7.91 2.75 2.07 13.96 2.21
  • No TCI is reported at 90% observation, because the remaining number of frames is insufficient to obtain a reliable estimate.